find the slope and the y intercept 6-1/4y=5x, the slope is??
Answer by ewatrrr(10682)


Hi


6-1/4y=5x OR 6-5x = y/4 OR 24 - 20x = y
y = -20x +24 |slope m = -20 and y-intercept is 24.
Line crossing the y axis at Pt(0,24), slanting to the left as m < 0
m = -20/1: as y increases/decreases by 20,
x increases/decreases by 1 in the opposite direction